OpenStudy (anonymous):

(x^1/2 -6^1/2)/x^1/2=0 y^3-(6)^3/2=0

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

is it simultaneous eqautions ? you just trying to solve for x and y ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you need a value of y and x or a function of y in terms of x ?

OpenStudy (anonymous):

value of x and y

OpenStudy (anonymous):

then we tackle them separately, y^3-6^3/2=0 then y^3=6^3/2 then y = (216/2)^(1/3) =4.76...... make sense ?